OpenStudy (anonymous):

The diameter of Z is 5 in. What is its area in terms of pi? (1 point) 2.5in.2 5in.2 6.25in.2 25in.2

OpenStudy (calculusxy):

Is there a certain figure that you need to find the area of ?

OpenStudy (anonymous):

Is Z a circular shape?

OpenStudy (anonymous):

yes

OpenStudy (calculusxy):

\[A = \pi r^2\]

OpenStudy (calculusxy):

Radius is always half of diameter 5/2 = 2.5

OpenStudy (calculusxy):

\[A = \pi (2.5)^2\]

OpenStudy (calculusxy):

We can have pi = 3.14 approximately

OpenStudy (anonymous):

so would it b c?

OpenStudy (anonymous):

would it be c?

OpenStudy (calculusxy):

Yes \[6.25\pi\]

OpenStudy (calculusxy):

We plugged in the pi since it said "in terms of pi"

OpenStudy (anonymous):

yes! thank you oh can you help me with another question?

OpenStudy (calculusxy):

Sure, I can try